OpenStudy (anonymous):

Quick Question: The slope-intercept form of a linear equation is y = mx + b, where x and y are coordinates of an ordered pair, m is the slope of the line, and b is where the line crosses the y-axis. Which is an equivalent equation solved for the slope, m? m = yx + b m = m = – b m = y –

OpenStudy (anonymous):

will and medal if you can solve

OpenStudy (anonymous):

Hand*

OpenStudy (anonymous):

alright first of all m=y2-y1/x2-x1

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

wait are those the exact choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually no hold up

OpenStudy (anonymous):

yx + b m = v-b/x m = v/x– b m = y – b/x

OpenStudy (anonymous):

First we subtract b from both sides, to get it out of the right side of the equation: y - b = mx + b - b y - b = mx Now we divide both sides by x, to get m completely by itself: (y-b)/x = (mx)/x (y-b)/x = m So m is equal to (y-b)/x.

OpenStudy (anonymous):

so you are simply solving for m in the equation

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

THE ANSWER IS M=